# 1st Siu-Wing Cheng    2nd Man Ting Wong

### v1

@@ I Introduction
In this paper, we adopt the linear measurement model in which $k$ measurements are bundled as a $k$-dimensional vector $b=Az$, where $A$ is a matrix in $\mathbb{R}^{k\times n}$. We are not concerned with the actual sensing mechanism that produces $b$ according to the linear measurement model. We are only concerned with the recovery of $z$ given $A$ and $b$.
Given $A$ and $b$, it has been proved that $z$ can recovered exactly by minimizing $\frac{1}{2}\|Ax-b\|^{2}+\eta\|x\|_{1}$ over all $x\in\mathbb{R}^{n}$ with an appropriate choice of $\eta>0$, provided that $A$ satisfies the restricted isometric property (RIP) of order $3s$ [3, 7, 11]. The matrix $A$ has order $3s$ if there exists $\delta\in(0,1)$ such that $(1-\delta)\|x\|^{2}\leq\|Ax\|^{2}\leq(1+\delta)\|x\|^{2}$ for all vectors $x\in\mathbb{R}^{n}$ with at most $3s$ nonzero coordinates. A popular way to construct a matrix $A\in\mathbb{R}^{k\times n}$ is to draw each entry of $A$ independently from $\mathcal{N}(0,1/k)$, i.e., each matrix entry is a Gaussian variable with mean zero and variance $1/k$. Such a matrix satisfies the RIP of order $s$ with probability at least $1-2e^{-\Theta(s)}$, provided that $k\geq Cs\ln(n/s)$ for some appropriate constant $C$ [3]. Random matrices are resilient to noise as well. On the other hand, it has been observed in the literature that it can be resource-demanding to solve $\min\frac{1}{2}\|Ax-b\|^{2}+\eta\|x\|_{1}$ or other similar optimization problems for large $n$. Also, there is no efficient algorithm to check whether a random matrix satisfies the RIP.
@@ I-B Related works
There has been a lot of work on constructing the sensing matrix $A$ deterministically that are based on codes, designs, algebraic curves, geometry over finite fields, and additive combinatorics (e.g. [2, 5, 6, 23, 24]). These matrices are usually sparse which leads to improved efficiency and facilitates the use of compressed sensing for larger problems. Nevertheless, such methods still resort to solving an optimization problem to recover the unknown signal.
@@
There is the related problem of 1-bit compressed sensing. In this setting, only the signs of the coordinates of $Az$ are given, and the problem is to recover an approximation of $z$ or just the support set of $z$. Binary iterative hard thresholding (BIHT) [19] is the widely celebrated algorithm for 1-bit compressed sensing. A closely related algorithm is the normalized BIHT (NBIHT) [16]. Various theoretical progresses have been obtained, including convergence of BIHT and NBIHT [29, 16], bounds for the sample size for 1-bit compressed sensing [1, 28], and computing the support set of $z$ or an approximation of $z$ in $o(n)$ time [39].
@@ III-A Setup
BIHT and NBIHT are designed for one-bit compressed sensing when only the signs of $Az$ are given. If a coordinate of $Az$ is postive, we convert it to +1+1; otherwise, we convert it to 1-1. One-bit compressed sensing may be a harder problem, but we include the comparisons with BIHT and NBIHT for completeness, as they can also determine the support set of $z$.
@@ III-B Accuracy
Let $x$ be the solution returned by a solver, and let $z$ be the ground truth of the unknown signal. To demonstrate the performance of support prediction for the algorithms, we define the ratio $R=|\mathrm{supp}(x)\cap\mathrm{supp}(z)|/|\mathrm{supp}(x)\cup\mathrm{supp}(z)|$ as the accuracy. This ratio equals $1$ when the predicted support is identical to the ground truth. It decreases when $\mathrm{supp}(x)$ does not contain all indices in $\mathrm{supp}(z)$ or when it includes false predictions outside $\mathrm{supp}(z)$. Across all methods, $R$ has an expected value at least 0.5 and a variance below $0.01$ over the trials. So $R$ is an effective accuracy indicator.

### v2

# 1st Man Ting Wong    2ndSiu-Wing Cheng

@@ I Introduction
In this paper, we adopt the linear measurement model in which $k$ measurements are bundled as a $k$-dimensional vector $b=Az$, where AA is a matrix in $\mathbb{R}^{k\times n}$. We are not concerned with the actual sensing mechanism that produces $b$ according to the linear measurement model. We are only concerned with the recovery of $z$ given AA and $b$.
Given AA and $b$, it has been proved that $z$ can recovered exactly by minimizing $\frac{1}{2}\|Ax-b\|^{2}+\eta\|x\|_{1}$ over all $x\in\mathbb{R}^{n}$ with an appropriate choice of $\eta>0$, provided that AA satisfies the restricted isometric property (RIP) of order $3s$ [3, 7, 11]. The matrix AA has order $3s$ if there exists $\delta\in(0,1)$ such that $(1-\delta)\|x\|^{2}\leq\|Ax\|^{2}\leq(1+\delta)\|x\|^{2}$ for all vectors $x\in\mathbb{R}^{n}$ with at most $3s$ nonzero coordinates. A popular way to construct a matrix $A\in\mathbb{R}^{k\times n}$ is to draw each entry of AA independently from $\mathcal{N}(0,1/k)$, i.e., each matrix entry is a Gaussian variable with mean zero and variance $1/k$. Such a matrix satisfies the RIP of order $s$ with probability at least $1-2e^{-\Theta(s)}$, provided that $k\geq Cs\ln(n/s)$ for some appropriate constant $C$ [3]. Random matrices are resilient to noise as well. On the other hand, it has been observed in the literature that it can be resource-demanding to solve $\min\frac{1}{2}\|Ax-b\|^{2}+\eta\|x\|_{1}$ or other similar optimization problems for large $n$. Also, there is no efficient algorithm to check whether a random matrix satisfies the RIP.
@@ I-B Related works
There has been a lot of work on constructing the sensing matrix AA deterministically that are based on codes, designs, algebraic curves, geometry over finite fields, and additive combinatorics (e.g. [2, 5, 6, 23, 24]). These matrices are usually sparse which leads to improved efficiency and facilitates the use of compressed sensing for larger problems. Nevertheless, such methods still resort to solving an optimization problem to recover the unknown signal.
@@
There is the related problem of 1-bit compressed sensing. In this setting, only the signs of the coordinates of AzAz are given, and the problem is to recover an approximation of $z$ or just the support set of $z$. Binary iterative hard thresholding (BIHT) [19] is the widely celebrated algorithm for 1-bit compressed sensing. A closely related algorithm is the normalized BIHT (NBIHT) [16]. Various theoretical progresses have been obtained, including convergence of BIHT and NBIHT [29, 16], bounds for the sample size for 1-bit compressed sensing [1, 28], and computing the support set of $z$ or an approximation of $z$ in $o(n)$ time [39].
@@ III-A Setup
BIHT and NBIHT are designed for one-bit compressed sensing when only the signs of AzAz are given. If a coordinate of AzAz is postive, we convert it to +1+1; otherwise, we convert it to 1-1. One-bit compressed sensing may be a harder problem, but we include the comparisons with BIHT and NBIHT for completeness, as they can also determine the support set of $z$.
@@ III-B Accuracy
Let $x$ be the solution returned by a solver, and let $z$ be the ground truth of the unknown signal. To demonstrate the performance of support prediction for the algorithms, we define the ratio $R=|\mathrm{supp}(x)\cap\mathrm{supp}(z)|/|\mathrm{supp}(x)\cup\mathrm{supp}(z)|$ as the accuracy. This ratio equals 11 when the predicted support is identical to the ground truth. It decreases when $\mathrm{supp}(x)$ does not contain all indices in $\mathrm{supp}(z)$ or when it includes false predictions outside $\mathrm{supp}(z)$. Across all methods, $R$ has an expected value at least 0.5 and a variance below $0.01$ over the trials. So $R$ is an effective accuracy indicator.